OpenStudy (anonymous):

Multiply: 6 times y times the square root of 7 • 2 the square root of y

OpenStudy (anonymous):

\[6y \times \sqrt{7}\times2\sqrt{y}=12y \sqrt{7y}\] Just multiply the stuff outside the square roots together and the stuff under the square roots togeter.

OpenStudy (cggurumanjunath):